# vid-TLDR: Training Free Token merging for Light-weight Video Transformer

### v1

@@ 4 Experiments
Video-text retrieval
First, we summarize the results of video-text retrieval with MSRVTT [68], MSVD [11], ActivityNet [9], and DiDeMo [2], LSMDC [49], Something-Something [30]. Video-text retrieval contains two subtasks: video-to-text retrieval, and text-to-video retrieval. Video-to-text retrieval is to find the most relevant text concerning the given video, while text-to-video retrieval is conducted in the opposite direction. We report the average of the results in Table 2. As summarized, our proposed method consistently shows competitive performances compared to base UMT [33] and outperforms ToMe[7] in every backbone and dataset. Compared to ToMe, vid-TLDR shows a performance gap of (+4.3%, +2.1%) R@1 in UMT-B, UMT-L even with the lower FLOPs on average. Further, vid-TLDR with UMT-B even surpasses the base model with the improvements of (+0.8%, +0.5%, +1.1%) R@1 while reducing FLOPs by (41.3%, 40.0%, 39.5%) in MSRVTT, MSVD, DiDeMo, respectively. And, we also observe similar improvements of (+0.2%, +2.3%) R@1 with UMT-L in MSVD and DiDeMo. We further provide the comparison with other video backbones in text-to-video retrieval (Table 3). For reporting the table, we experiment with the model used in Table 2. Although we largely reduce the computational cost of UMT-B, and UMT-L by 36.5%, 42.7% on average, they still show superior performance compared to other video backbones in MSRVTT, DiDeMo, and ActivityNet.
Video question answering.
We experiment with video question answering with MSR-QA [67] and MSVD-QA [67], summarizing the results in Table 4. In MSR-QA, the official number of UMT-B and UMT-L is 44.9% and 47.1%, respectively. After adopting vid-TLDR on each model, we achieved the competitive performance of 44.8% and 47.0% while reducing FLOPs by 37.9% in UMT-B, and 42.1% in UMT-L. Similarly, we could lessen the computational cost in MSVD-QA with the small performance degradation despite the much lower FLOPs. Specifically, the performance drop is only 0.1% and 0.3% in UMT-B and UMT-L with much lower FLOPs compared to the original model. To summarize, vid-TLDR boosts the model with a minor accuracy drop in video question answering as well as video-text retrieval.
@@
vid-TLDR with other video Transformers and tasks.
To demonstrate the generalizability of vid-TLDR with other video Transformers and tasks, we apply vid-TLDR to other video Transformers (VideoMAE [51], and ViFi-CLIP [48]. The results of action recognition with VideoMAE are in Table 5, and the result of base-to-novel generalization (i.e., training only with the base classes, then evaluating on both seen (base) and unseen (novel) classes) with ViFi-CLIP are in Table 6. We use Something Something V2 [19] for action recognition and UCF101 [50] for both action recognition and base-to-novel generalization tasks. Experimental results are promising that vid-TLDR lessens the complexity of VideoMAE by almost 70% with a minimal performance degradation compared to ToMe (e.g., 14.4% improvement over ToMe in UCF101). Also, in the base-to-novel generalization, vid-TLDR halving the FLOPs of ViFi-CLIP surpassing ToMe with a 19.4% gain in the harmonic mean (HM).
@@ 4.2 Ablation studies
In this section, we provide the ablations studies of vid-TLDR. We studied the effectiveness of each proposed component with UMT-B [33] and MSRVTT [68] and show the results in Table 7. The first row of the table indicates the base model without any token reduction. First, regarding the metric for informativeness, we have compared our proposed saliency scores $s$ in Equation 9 with the attentiveness $\bar{a}$ in Equation 3, and attention rollout $\tilde{a}$ in Equation 4. For comparison, based on each metric, we simply prune the token having the low scores, equal to the number used for reporting Table 2. As shown in the table, since the tokens are dropped in the earlier layers, the attentiveness and attention rollout may drop the salient tokens resulting in a substantial performance drop in both text-to-video retrieval and video-to-text retrieval. Specifically, the performance degradation on average is 1.0%, 0.7% when using $\bar{a}$ and $\tilde{a}$. Further, although attention rollout shows slightly better performance than simple attentiveness, it shows worse FLOPS due to the repeated forward process for estimating attention flows. On the other hand, by dropping the tokens based on our saliency scores $s$ estimated by the sharpness of attention, we could lessen the FLOPs with superior performance compared to base. Finally, introducing the saliency-aware token merging with background drop and foreground sharpening, we have achieved the +0.8% gain despite the 58.7% FLOPs compared to the original UMT.
 $\bar{a}$ $\tilde{a}$ $s$ B.D & F.S GFLOPS T2V V2T Mean
@@
Table 7: Ablations studies on vid-TLDR. The first row indicates the base UMT-B [33] without vid-TLDR. B.D. and F.S. denote the background drop and foreground sharpening in saliency-aware token merging.
@@ 4.3 Analysis
Visualization of vid-TLDR results.
In Figure 6, we show the qualitative results to understand the behavior of vid-TLDR. Given video clips of MSRVTT, we visualize the merged token. We further provide the heatmaps for analyzing the mass of each token. More precisely, we divide the mass of merged tokens by the number of constituent tokens to represent the mass concerning each input token. As shown in the figure, through the saliency-aware token merging, the foreground object shows a much higher mass than the background tokens. In short, through Equation 7, our vid-TLDR minimizes the hindrance from the background tokens during the self-attention layer.
@@ 5 Related works
Vision Transformer.
ViT [14] has become one of the most popular and basic components in the computer vision domain along with CNNs. With the surge of large-scale datasets, its lower inductive bias compared to CNNs endows it with robust generalization, leading to another successive adoption on several downstream tasks in computer vision, including classification [14, 52], detection [10, 61, 43, 77, 32, 75] segmentation [59, 64], image encoding [21, 3, 62] for generation [28]. Although Transformer architecture shows promising results on computer vision the attention mechanism [4] incurs quadratic complexity and restrains its scalability. So, there have been many attempts to mitigate this problem. These attempts can be categorized as ameliorating model architecture itself, leveraging pre-trained models, or reducing the input tokens. For example, [23, 54, 58, 25, 74] focused on the attention mechanism itself to approximate the complexity to linear. More recent works enabled acceleration even without model modification by pruning [44, 24, 27] or merging [8, 7, 42, 45] input tokens, with minimal performance degradation.
Video understanding.
Video understanding is not the same as the image, since frames of video are not independent images, but highly related to each other in the temporal axis. Prior works have leveraged transformers for understanding video, including retrieval [16, 30, 55, 12, 40, 70, 33, 56], question answering [31, 71, 56, 76, 30, 12, 55, 60, 70], captioning [37, 20] and representation learning [51, 57, 22]. Especially, along with the success of image foundation models, some works [48, 41] leverage CLIP [46], which is pre-trained ViTs with the large-scale image-text pairs, for understanding video. Also, recent studies [76, 51, 57, 33, 60] focus on scaling the Transformers for video foundation models to utilize the flexibility for multi-modal tasks. Yet, despite the intensive computational cost caused by the massive number of tokens, efficient video Transformers are less explored.
@@ Appendix B Detailed reduced number of tokens
Appendix C Zero-shot retrieval results
To validate the effectiveness of vid-TLDR, we report additional zero-shot retrieval results for baseline model [33], existing token merging method (ToMe [8]) and ours across MSVD [11], MSRVTT [68], DiDeMo [2], ActivityNet [9]. All experiments were conducted with evaluation only, without additional training. The average results for text-to-video and video-to-text retrieval accuracy across these four datasets are presented in Table B, C, D, E, and F. We conducted a grid search to determine the effective number of tokens to reduce. In the case of ToMe, tokens were merged uniformly across layers following the default configuration, and a grid search was conducted with 20-25 tokens per layer. On the other hand, our method merged in the first four front layers, exploring total reductions of 300, 330, and 360 tokens. Across all configurations, vid-TLDR consistently achieves higher R@1 with lower FLOPs compared to ToMe. On average, it achieves 25.0% and 11.8% reduction in FLOPs, along with 2.6% and 2.2% higher R@1 for UMT-B and UMT-L.
Table B: Comparison of zero-shot retrieval performance for four datasets.
@@ Appendix C Zero-shot retrieval results
Table C: Detailed zero-shot retrieval performance on MSVD.
@@
Table D: Detailed zero-shot retrieval performance on MSRVTT.
@@
Table E: Detailed zero-shot retrieval performance on DiDeMo
@@
Table F: Detailed zero-shot retrieval performance on Activitynet.

### v2

@@ 4 Experiments
Video-text retrieval First, we summarize the results of video-text retrieval with MSRVTT [68], MSVD [11], ActivityNet [9], and DiDeMo [2], LSMDC [49], Something-Something [30]. Video-text retrieval contains two subtasks: video-to-text retrieval, and text-to-video retrieval. Video-to-text retrieval is to find the most relevant text concerning the given video, while text-to-video retrieval is conducted in the opposite direction. We report the average of the results in Table 2. As summarized, our proposed method consistently shows competitive performances compared to base UMT [33] and outperforms ToMe[7] in every backbone and dataset. Compared to ToMe, vid-TLDR shows a performance gap of (+4.0%, +2.1%) on average R@1 in UMT-B, UMT-L even with the lower FLOPs. Further, vid-TLDR with UMT-B even surpasses the base model with the improvements of (+0.8%, +0.6%, +0.3%) R@1 while reducing FLOPs by (41.3%, 40.2%, 29.8%) in MSRVTT, MSVD, DiDeMo, respectively. And, we also observe that vid-TLDR achieves competitive performance with base UMT-L despite the much lower FLOPs. We further provide the comparison with other video backbones in text-to-video retrieval (Table 3). For reporting the table, we experiment with the model used in Table 2. Although we largely reduce the computational cost of UMT-B, and UMT-L by 34.1%, 42.7% on average, they still show superior performance compared to other video backbones in MSRVTT, DiDeMo, and ActivityNet.
Video question answering. We experiment with video question answering with MSR-QA [67] and MSVD-QA [67], summarizing the results in Table 4. In MSR-QA, the results of UMT-B and UMT-L are 44.9% and 47.1%, respectively. After adopting vid-TLDR on each model, we achieved the competitive performance of 44.8% and 47.0% while reducing FLOPs by 37.9% in UMT-B, and 42.1% in UMT-L. Similarly, we could lessen the computational cost in MSVD-QA with the small performance degradation despite the much lower FLOPs. Specifically, the performance drop is only 0.1% and 0.3% in UMT-B and UMT-L with much lower FLOPs compared to the original model. To summarize, vid-TLDR boosts the model with a minor accuracy drop in video question answering as well as video-text retrieval.
@@
vid-TLDR with other video Transformers and tasks. To demonstrate the generalizability of vid-TLDR, we apply vid-TLDR to other video Transformers: VideoMAE [51], and ViFi-CLIP [48]. The results of action recognition with VideoMAE are presented in Table 5. We use Something Something V2 (SSV2[19] and UCF101 [50]. Experimental results are promising that vid-TLDR lessens the complexity of VideoMAE by almost 70% (180.5 (G) $\rightarrow$ 56.1 (G)) with a minimal performance degradation compared to ToMe (e.g., 14.4% (75.7 $\rightarrow$ 90.1) improvement over ToMe). Further, we adopt vid-TLDR on ViFi-CLIP in the base-to-novel generalization tasks (i.e., training only with the base classes, then evaluating on both seen (base). As shown in Table 6 with UCF-101, vid-TLDR halving the FLOPs (563 $\rightarrow$ 279) of ViFi-CLIP surpassing ToMe with a 19.4% (55.3% $\rightarrow$ 74.7%) gain in the harmonic mean (HM).
@@ 4.2 Ablation studies
In this section, we provide the ablations studies of vid-TLDR. We studied the effectiveness of each component with UMT-B [33] and MSRVTT [68] (Table 7). The first row of the table indicates the base model without any token reduction. First, regarding the metric for informativeness, we have compared our proposed saliency scores $s$ in Equation 9 with the attentiveness $\bar{a}$ in Equation 3, and attention rollout $\tilde{a}$ in Equation 4. For comparison, based on each metric, we simply prune the token having the low scores, equal to the number used for reporting Table 2. As shown in the table, since the tokens are dropped in the earlier layers, the attentiveness and attention rollout may drop the salient tokens resulting in a substantial performance drop in both text-to-video retrieval and video-to-text retrieval. Specifically, the performance degradation on average is 1.0%, 0.7% when using $\bar{a}$ and $\tilde{a}$. Further, although attention rollout shows slightly better performance than simple attentiveness, it shows worse FLOPS due to the repeated forward process for estimating attention flows. On the other hand, by dropping the tokens based on our saliency scores $s$ estimated by the sharpness of attention, we could lessen the FLOPs with superior performance compared to base. Finally, introducing the saliency-aware token merging, we have achieved the +0.8% (50.0% $\rightarrow$ 50.8%) gain despite the 58.7% FLOPs compared to the original UMT.
 $\bar{a}$ $\tilde{a}$ $s$ S.A. ToMe GFLOPS T2V V2T Mean
@@
Table 7: Ablations studies on vid-TLDR. The first row indicates the base UMT-B [33]. S.A. ToMe denotes the saliency-aware token merging.
@@ 4.3 Analysis
Visualization of vid-TLDR results. In Figure 6, we show the qualitative results to understand the behavior of vid-TLDR. Given video clips of MSRVTT, we visualize the merged token. We further provide the heatmaps for analyzing the mass of each token. More precisely, we divide the mass of merged tokens by the number of constituent tokens to represent the mass concerning each input token. As shown in the figure, through the saliency-aware token merging, the foreground object shows a much higher mass than the background tokens. In short, through Equation 7, our vid-TLDR minimizes the hindrance from the background tokens during the self-attention layer.
@@ 5 Related works
Vision Transformer. ViT [14] has become one of the most popular and basic components in computer vision along with CNNs. With the surge of large-scale datasets, its lower inductive bias compared to CNNs endows it with robust generalization, leading to another successive adoption on several downstream tasks in computer vision, including classification [14, 52], detection [10, 61, 43, 77, 32, 75] segmentation [59, 64], image encoding [21, 3, 62] for generation [28]. Although Transformer shows promising results on computer vision the attention mechanism [4] incurs quadratic complexity and restrains its scalability. So, there have been many attempts to mitigate this problem. These attempts can be categorized as ameliorating model architecture itself, leveraging pre-trained models, or reducing the input tokens. For example, [23, 54, 58, 25, 74] focused on the attention mechanism itself to approximate the complexity to linear. More recent works enabled acceleration even without model modification by pruning [44, 24, 27] or merging [8, 7, 42, 45] input tokens, with minimal performance degradation.
Video understanding. Video understanding is not the same as the image, since frames of video are not independent images, but highly related to each other in the temporal axis. Prior works have leveraged transformers for understanding video, including retrieval [16, 30, 55, 12, 40, 70, 33, 56], question answering [31, 71, 56, 76, 30, 12, 55, 60, 70], captioning [37, 20] and representation learning [51, 57, 22]. Especially, along with the success of image foundation models, some works [48, 41] leverage CLIP [46], which is pre-trained ViTs with the large-scale image-text pairs, for understanding video. Also, recent studies [76, 51, 57, 33, 60] focus on scaling the Transformers for video foundation models to utilize the flexibility for multi-modal tasks. Yet, despite the intensive computational cost caused by the massive number of tokens, efficient video Transformers are less explored.
@@ 6 Conclusion
Acknowledgments
This work was partly supported by ICT Creative Consilience Program through the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT)(IITP-2024-2020-0-01819), the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT)(NRF-2023R1A2C2005373), and Electronics and Telecommunications Research Institute (ETRI) grant funded by the Korean government (24ZB1200, Fundamental Technology Research for Human-Centric Autonomous Intelligent Systems).
@@ Appendix B Detailed reduced number of tokens
Appendix C Detailed video-text retrieval
In Table 2 of the main paper, we provide video-text retrieval by averaging the performances of video-to-text retrieval and text-to-retrieval. Here, we provide the detailed results of each task in Table B with MSRVTT [68], MSVD [11], ActivityNet [9], DiDeMo [2], LSMDC [49], SSV2-Label/Template [30].
Table B: Detailed results on text-to-video & video-to-text retrieval.
    Dataset Metric Type         UMT-B         UMT-L
Base ToMe Ours Base ToMe Ours
MSRVTT GFLOPs $\downarrow$ (G) 303.3 231.4 178.0 984.6 529.7 563.1
R@1 $\uparrow$ T2V 51.0 46.6 50.9 58.8 55.6 58.1
V2T 49.0 47.4 50.7 58.6 56.0 58.7
R@5 $\uparrow$ T2V 76.5 71.8 75.8 81.0 79.7 81.0
V2T 77.0 74.6 75.5 81.6 79.4 81.6
R@10 $\uparrow$ T2V 84.2 80.9 83.9 87.1 85.3 86.8
V2T 84.7 83.2 83.9 86.5 86.8 86.9
MSVD GFLOPs $\downarrow$ (G) 303.3 218.7 181.3 984.6 574.5 563.1
R@1 $\uparrow$ T2V 50.8 47.8 50.5 58.2 57.1 57.9
V2T 73.3 71.3 74.9 82.4 81.9 82.7
R@5 $\uparrow$ T2V 79.7 78.2 79.7 83.9 83.6 83.8
V2T 88.8 89.4 90.0 94.6 93.9 94.5
R@10 $\uparrow$ T2V 85.7 85.5 86.3 89.6 89.4 89.4
V2T 93.7 92.5 93.3 96.7 96.0 96.3
ActivityNet GFLOPs $\downarrow$ (G) 303.3 236.8 227.6 984.6 574.5 572.9
R@1 $\uparrow$ T2V 58.3 51.9 57.8 66.8 63.3 66.7
V2T 56.0 51.4 55.4 64.4 61.7 63.9
R@5 $\uparrow$ T2V 83.9 80.8 83.7 89.1 87.3 88.6
V2T 83.5 80.6 83.2 89.1 86.5 88.7
R@10 $\uparrow$ T2V 91.5 89.4 91.5 94.9 93.6 94.4
V2T 91.7 89.9 91.1 94.8 93.6 94.5
DiDeMo GFLOPs $\downarrow$ (G) 303.3 218.7 212.8 984.6 574.5 559.0
R@1 $\uparrow$ T2V 63.7 56.4 64.1 72.5 68.7 72.3
V2T 60.5 58.2 60.8 69.2 67.3 68.5
R@5 $\uparrow$ T2V 87.8 82.3 87.2 91.2 89.5 91.2
V2T 85.7 82.8 85.1 89.9 89.2 89.8
R@10 $\uparrow$ T2V 93.5 89.0 92.7 94.8 94.6 94.2
V2T 90.7 89.5 90.4 94.1 92.9 93.8
LSMDC GFLOPs $\downarrow$ (G) 303.3 223.2 206.2 984.6 574.5 583.7
R@1 $\uparrow$ T2V 32.7 25.6 32.9 43.0 39.7 43.1
V2T 32.7 29.0 31.8 41.4 38.7 40.7
R@5 $\uparrow$ T2V 54.7 47.6 53.6 65.5 61.8 64.5
V2T 53.5 50.5 52.9 64.3 61.4 63.6
R@10 $\uparrow$ T2V 63.4 57.6 63.3 73.0 61.8 71.4
V2T 63.2 57.0 63.1 71.5 67.9 70.2
SSV2-label GFLOPs $\downarrow$ (G) 303.3 232.2 212.9 984.6 627.2 610.9
R@1 $\uparrow$ T2V 64.1 58.7 64.3 73.3 69.4 73.1
V2T 63.8 61.7 63.3 71.5 70.3 71.1
R@5 $\uparrow$ T2V 88.2 84.9 88.0 93.7 92.1 93.3
V2T 88.3 87.2 87.4 93.1 92.3 92.7
R@10 $\uparrow$ T2V 92.7 90.8 92.6 96.6 95.8 96.6
V2T 93.2 92.4 92.9 96.7 95.8 96.4
SSV2-Template GFLOPs $\downarrow$ (G) 303.3 241.4 203.7 984.6 627.2 572.9
R@1 $\uparrow$ T2V 87.9 85.1 87.4 90.8 88.5 90.2
V2T 61.3 59.5 60.6 66.1 65.5 66.0
R@5 $\uparrow$ T2V 99.4 100.0 99.4 99.4 98.9 100.0
V2T 88.5 87.9 87.4 92.3 91.4 91.7
R@10 $\uparrow$ T2V 100.0 100.0 100.0 100.0 100.0 100.0
V2T 93.7 93.1 92.6 95.6 95.5 95.8
Appendix D Zero-shot retrieval results
To validate the effectiveness of vid-TLDR, we report additional zero-shot retrieval results for baseline model [33], existing token merging method (ToMe [8]) and ours across MSVD [11], MSRVTT [68], DiDeMo [2], ActivityNet [9]. All experiments were conducted with evaluation only, without additional training. The average results for text-to-video and video-to-text retrieval accuracy across these four datasets are presented in Table C, D, E, F, and G. We conducted a grid search to determine the effective number of tokens to reduce. In the case of ToMe, tokens were merged uniformly across layers following the default configuration, and a grid search was conducted with 20-25 tokens per layer. On the other hand, our method merged in the first four front layers, exploring total reductions of 300, 330, and 360 tokens. Across all configurations, vid-TLDR consistently achieves higher R@1 with lower FLOPs compared to ToMe. On average, it achieves 25.0% and 11.8% reduction in FLOPs, along with 2.6% and 2.2% higher R@1 for UMT-B and UMT-L.
Table C: Comparison of zero-shot retrieval performance for four datasets.
@@ Appendix D Zero-shot retrieval results
Table D: Detailed zero-shot retrieval performance on MSVD.
@@
Table E: Detailed zero-shot retrieval performance on MSRVTT.
@@
Table F: Detailed zero-shot retrieval performance on DiDeMo
@@
Table G: Detailed zero-shot retrieval performance on Activitynet.